OpenStudy (anonymous):

If f(x)=sin^3 (4x), find f'(pi/3).

OpenStudy (jennychan12):

take f'(x). \[f'(x) = 3(\sin(4x))^2(4)\] \[f'(x) = 12(\sin(4x))^2)\] plug in \[\frac{ \pi }{ 3 }\] and you have the answer.

OpenStudy (anonymous):

A. -3/2, B. -9/2, C. -3sqrt(3), D. 3sqrt(3), E. -9/8. Which is the answer? When I plug in the value that you said, none of them are correct.

OpenStudy (jennychan12):

shoot. i did f'(x) wrong. sorry. \[f'(x) = 3(\sin(4x))^2(\cos(4x))(4)\] simplify. then plug in.

OpenStudy (jennychan12):

the answer is one of the choices. i'll let you figure it out,

OpenStudy (anonymous):

hint when you plug in \[\frac{ \pi }{ 3}\] to sin(4x).....what is \[\sin(\frac{ 4\pi }{ 3 })\]

OpenStudy (anonymous):

Thanks, guys.